OpenStudy (anonymous):

i need each of this number as a power: 25, 64, 8, 27

OpenStudy (anonymous):

5^2, 8^2, 2^3, 3^3

OpenStudy (anonymous):

25=5^2 64=8^2 8=2^3 27=3^3